OpenStudy (anonymous):

Solve for x: 2x^2 - 4x - 14 = 0

OpenStudy (anonymous):

\[x^2 -2x - 7=0\] \[x^2 - 2x - 7 +1 -1 =0\] \[ (x-1)^2= 8\] \[x -1=\pm 2\sqrt{2}\] \[x = 1 \pm 2\sqrt{2}\]